# Pseudospectra of Quasinormal Modes and Holography

### v1

@@ 4 Discussion
We note that, as we have emphasized, the definition of the holographic Green’s function implicitly relies on analytic continuation across the horizon. This analyticity requirement is much stronger than the requirement of existence of the energy norm. A way to circumvent this has been suggested in (26) and consists in replacing the energy norm by a Sobolev norm. In physicists terms this corresponds to higher order derivative terms in the norm. Higher derivative terms up to $|\partial_{\rho}^{n}a|^{2}$ amount to lowering the limit for integrability to $\Im(\omega)<1-2n$. In order to recover the exact spectrum one would of course have to take a limit with infinitely many derivatives. From the physics point of view the significance of such higher derivative terms is not clear.
@@
Another line of thought could be that one considers the underlying theory (being it a scalar field, a Maxwell field or the metric itself) as an effective field theory valid down to a finite cutoff length scale $\Lambda$. Then we would necessarily have some huge but finite value for $N$ determined e.g. by the criterion that the minimal distance between points of the discretization is larger than $\Lambda$. Alternatively, one could also impose the boundary conditions not directly at the horizon but slightly outside at a sort of “stretched” horizon (29).
@@
Finally let us close this overview by pointing to additional results on quasinormal modes in Anti-de Sitter space. The pseudospectrum in infalling Eddingtion-Finkelstein coordinates has been investigated in (30). One of the main findings was that in certain cases the pseudospectrum can significantly reach up into the upper halfplane giving rise to possible transient behaviour. Structural aspects of the pseudospectrum of quasinormal modes for AdS black holes have been pointed out and further investigated in (28). In particular the results in infalling Eddington-Finkelstein and regular coordinates have been contrasted. The properties of the pseudospectrum of black hole metrics have also been shown to give rise to transient behaviour for which a sum of $M$ quasinormal modes can be long lived of order $\log(M)$ in (31).
@@ Acknowledgments
We thank V. Boyanov and especially J. Jaramillo for numerous insightful discussions on the properties of quasinormal modes and their pseudospectra.
Conflict of Interest Statement
The authors declare that the research was conducted in the absence of any commercial or financial relationships that could be construed as a potential conflict of interest.
Author Contributions
All authors contributed equally.

### v2

@@ III Discussion
We note that, as we have emphasized, the definition of the holographic Green’s function implicitly relies on analytic continuation across the horizon. This analyticity requirement is much stronger than the requirement of existence of the energy norm. A way to circumvent this has been suggested in Warnick (2015) and consists in replacing the energy norm by a Sobolev norm. In physicists terms this corresponds to higher order derivative terms in the norm. Higher derivative terms up to $|\partial_{\rho}^{n}a|^{2}$ amount to lowering the limit for integrability to $\mathop{\rm Im}(\omega)<1-2n$. In order to recover the exact spectrum one would of course have to take a limit with infinitely many derivatives. From the physics point of view the significance of such higher derivative terms is not clear.
@@
Another line of thought could be that one considers the underlying theory (being it a scalar field, a Maxwell field or the metric itself) as an effective field theory valid down to a finite cutoff length scale $\Lambda$. Then we would necessarily have some huge but finite value for $N$ determined e.g. by the criterion that the minimal distance between points of the discretization is larger than $\Lambda$. Alternatively, one could also impose the boundary conditions not directly at the horizon but slightly outside at a sort of “stretched” horizon (Price and Thorne, 1986).
Let us now emphasize the importance of pseudospectra in the context of holography. From the gravitational side, pseudospectra probe how much the quasinormal frequencies change if the background is slightly modified in some way (e.g., by a change of the geometry and/or the background value of the fields). Consequently, in the dual quantum field theory, pseudospectra help us estimate how much the poles of the retarded Green’s functions might change if the theory is slightly perturbed. In both cases, these perturbations should be understood as perturbations to the Lagrangian, leading to a change in the spectrum of excitations. Then, spectral instability suggests that holographic models might not be able to accurately capture the actual spectra of real physical systems such as quark-gluon plasma. However, valuable information, such as transient dynamics, can still be obtained by studying pseudospectra Jaramillo (2022); Trefethen and Embree (2005).
@@
We shall now point to additional results on quasinormal modes in Anti-de Sitter space. The pseudospectrum in infalling Eddingtion-Finkelstein coordinates has been investigated in (Cownden et al., 2024). One of the main findings was that in certain cases the pseudospectrum can significantly reach up into the upper halfplane giving rise to possible transient behaviour. Structural aspects of the pseudospectrum of quasinormal modes for AdS black holes have been pointed out and further investigated in Boyanov et al. (2024). In particular the results in infalling Eddington-Finkelstein and regular coordinates have been contrasted. The dependence of pseudospectra on the choice of coordinates still needs further investigation. The properties of the pseudospectrum of black hole metrics have also been shown to give rise to transient behaviour for which a sum of $M$ quasinormal modes can be long lived of order $\log(M)$ in Carballo and Withers (2024). The stability of complex linear momenta ($\mathbb{C}$LMs) in Anti-de Sitter was studied in Garcia-Fariña et al. (2024). Remarkably the pseudospectrum of $\mathbb{C}$LMs was observed to be convergent allowing for quantitative results.
In this paper we have reviewed the holographic perspective on quasinormal modes and quasinormal frequencies of AdS black holes. In this context, pseudospectrum analysis offers an invaluable tool for assessing the stability as well as investigating the existence of transient dynamics. Numerically-computed pseudospectra do not converge in the energy norm because outgoing modes can still have finite energy. We believe that the lack of convergence is not a flaw of the construction but rather a fundamental feature which needs to be addressed using a physics-motivated regulator.
@@ Acknowledgments
We thank V. Boyanov and especially J.L. Jaramillo for numerous insightful discussions on the properties of quasinormal modes and their pseudospectra.